7. Solve each of the following exponential equations. x-2 a. 42*+1 64
Answer #1

a) 4^(2x+1) = 64

4^(2x+1) = 4^3

2x +1 = 3

2x = 2

x = 1

b) 3^(x-2)/3 = sqrt 27

3^(x-2)/3 = 27^1/2

3^(x-2)/3 = 3^3/2

( x- 2 ) / 3 = 3/2

cross multiplying

2(x-2) = 9

2x - 4 = 9

2x = 13

x = 13/2